# Photoinduced Charge Transfer in Transition Metal Dichalcogenide Quantum Dots

### v1

@@ Abstract
In this paper, a charge transfer mechanism in transition metal dichalcogenide (TMDC) quantum dots (QDs) is studied. A common dye, Rhodamine 6G (R6G), has been used as fluorescent molecule and QDs of molybdenum disulfide () and tungsten disulfide () have been used as acceptors of electrons in the photo-induced charge transfer process. The MoS2 and WS2 QDs have been synthesized using ultrasonication method. These QDs have been characterized using transmission electron microscope (TEM), UV-Vis spectrophotometer, and fluorimeter. The TEM images of these QDs show well dispersed particles of size 2 nm. These QDs show intense fluorescence emission when excited with light having wavelength, 350 nm. In presence of light, the photons help in creating charges in fluorescent dye molecule, and TMDCs QDs felicitate the charge transfer in the process. The charge transfer phenomenon has been studied using time correlated single photon counting (TCSPC) set-up, which is a time resolved fluorescence spectroscopic technique. The time resolved fluorescence study shows a drastic change in the fluorescence (FL) lifetime of R6G molecules in the presence of QDs. This decrease in FL lifetime of R6G shows that the MoS2, and WS2 QDs provide additional path to the photo-generated electrons in the excited state of R6G. This study can be further extended to optoelectronic devices, where charge transfer plays an important role in the device efficiency and performance.
@@
In recent years, various types of quantum dots (QDs) have been widely explored due to their vast applications in the field of cell imaging, photovoltaic devices, light emitting devices, etc. Xu et al. (2021); Li et al. (2024); Wang et al. (2020); Sadeghi et al. (2019); Li et al. (2016); Xue et al. (2019); Meng et al. (2024) The most attractive property of these materials are being size dependent opto-electronic properties, which help these materials to be a suitable candidate for designing various opto-electronic devices.Li et al. (2015); Litvin et al. (2017) The intriguing properties of QDs have been explored in a composite system or device through their composites with other materials, which have been explored to various molecules, QDs, 2D nanomaterials, etc. When the QDs are in composite form with biomolecules such materials have been used for bio-imaging. While, the core-shell configuration of QDs have been explored for optoelectronic devices. These QDs have been sensitized with other semiconductors and have created inexpensive, next-generation photovoltaic devices for various applications. In these applications, QDs have been favored because of their size-dependent optical and electronic properties. In this regard, interaction of QDs with surround media or materials is critical for device performance. Although electronic interactions between organic molecules and fluorescent QDs have been well explored, the system where QDs are coupled to other inorganic molecules, such paired system have been elucidated to a lesser extent. These kinds of interactions are different from those in QDs- molecular systems because the inorganic material possesses a continuum electronic states, while the molecular acceptors have discrete electronic states. When the QDs are coupled with other inorganic materials, the QDs can be used to donate electrons, which have been observed in photovoltaic devices, or these QDs can act as donor/acceptor of the electrons, which have been explored in light emitting diodes (LEDs). In the applications where these QDs are used into devices, the electron transfer is involved with a proper direction of the transfer. As resultant, to understand the factors which use to drive the electron transfer in these composite systems is critical for better understanding and further exploitation of the unique properties of QDs in optoelectronic devices.Algar and Krull (2008); Wang et al. (2021); Zhu, Zhang, and Xia (2018); Lima et al. (2022); Wu et al. (2020); Rani et al. (2020)
@@
There are various studies on different kind of QDs, where these QDs have been used as acceptor/donor or both of the electrons. They have also been explored for donor or acceptor in the resonance energy transfer. In recent years researchers have explored various kinds of layered based QDs in place of traditional QDs, e.g., CdS, CdSe, PbS, etc.Kharangarh et al. (2023); Sariga et al. (2023); Gan et al. (2015); Guo et al. (2019) In optoelectronic-based applications these QDs act as light harvesting and charge separation centers, and an important step is the electron transfer or hole transfer from the photoexcited QDs to electron or hole acceptors near the QDs through a nanoscale interface. The overall device efficiency depends on these charge transfer processes with competing loss pathways. Therefore, understanding and controlling interfacial charge transfer processes is essential for the improvement of QDs based devices. The size of the QDs also plays an important role in the charge transfer process. It has been observed that there is increase in electron transfer rate with decreasing QD size. Zhu et al. (2016)
@@
In recent years, various 2D layered nanomaterials have been turned in to QDs, and these layered QDs exhibit quite different properties as compared to their 2D structures and as well as bulk counterparts. In this regards, 2D layered structure of transition metal dichalcogenide (TMDC) have also been converted to QDs using various methods. In these TMDC based QDs, MoS2, and WS2 QDs have been studied extensively. They possess interesting optical and electronic properties as their electronic structure can be tuned by tunning their size and surface morphology. These QDs have been used in various applications like sensors, LEDs, photodetectors, cell imaging, etc. Yan et al. (2016); Ghorai et al. (2017); Singh et al. (2019); Guo and Li (2020)
@@
There are a few studies where TMDC QDs have been used as donor of energy in various hybrid systems, due to their photo-luminescence in the UV region. There are numerous studies on the energy transfer/ charge transfer in different set of nanomaterials and there applications in sensors, photodetectors, solar cells, etc. In this work, we have studied the charge transfer in two different kinds of fluorescent materials; fluorescent dye molecules, and QDs. Conventionally, QDs are used as donor/acceptor and the dye molecules as acceptor/donor in the energy transfer, and most of the studies where dye-QDs are used are centered around the energy transfer and its various applications. Here, we have explored the possibility of charge transfer from dye molecules to QDs, which would provide an important insight while developing optoelectronic devices based on fluorescent nanomaterials. In this paper, we have used a common dye molecule; R6G as source of photo-excited electrons, and TMDCs QDs as acceptors for the electrons.
@@
The TMDC QDs were synthesized using top-down strategy as mentioned in Long et al. with slight modification.Long et al. (2016) In typical synthesis procedure, 0.05 g of TMDC powder was dissolved in 50 mL of NMP solution with constant stirring for 15 hours at 29 ∘C. The resulting dark suspension was centrifuged at 10,000 rpm for 40 min. The supernatant of the solution was collected and used for further studies.
@@
In our study, prior to exploring charge transfer mechanism, these QDs were characterized by transmission electron microscopy (TEM), fluorescence spectroscopy, and absorption spectroscopy. The TEM images of these QDS were recorded using JEOL-2010 TEM setup operated at 200 kV. The absorption spectra of the QDs were recorded using Shimadzu (UV-2450) UV-Vis spectrophotometer. In order to study charge transfer in these TMDCs QDS, the QDs were mixed with a fixed quantity of R6G solution. For charge transfer studies, 0.5 mL aqueous solution of 10-6 M R6G was further diluted to 3.5 mL, and 1 mL of QDs solution was mixed to R6G solution. In all the samples the NMP and water contents were kept same. The steady state fluorescence spectra of R6G molecules were recorded in absence and presence of TMDCs QDs by using fluorimeter with excitation wavelength of 430 nm. The time resolved fluorescence study was performed using time-correlated single-photon counting module (TCSPC) (FL920, Edinburgh Instruments, UK) with excitation at 465 nm with diode laser as the excitation source.
The absorption spectra of MoS2 and WS2 quantum dots are shown in Fig 1.
@@
A clear absorption peak around 325 nm is observed, which is a characteristic peak of WS2 quantum dots and it is due to excitonic features of the QDs. Similarly, for MoS2 QDs a band edge around 380 nm is observed. The color of suspenssion and the absorption peaks are similar to the previous reported results. Coleman et al. (2011); Long et al. (2016) When the MoS2 and WS2 are in the bulk or nanosheet form, they exhibit quite different absorption patterns having two distinct peaks around 600 nm which are known as and excitonic peaks. While in case of QDs these and excitonic peaks disappear and absorption is observed in UV region. The strong blue shift in the absorption of the as-prepared MoS2 and WS2 QDs is attributed to the quantum confinement effect and edge effects when the lateral size of the QDs is reduced to few tens of nanometer. Li et al. (2017); Liu et al. (2014) In order to have insight of the band gaps of these QDs, the Tauc plots of these QDs are drawn. The Tauc plots of the quantum dots are shown in Fig 2.
@@
In case of semiconductors, it has been observed the bandgap increases with decrease in the size of the semiconducting quantum dots. Wu et al. (2005) The size dependent bandgap is given by following equation:
@@
where is the bandgap of the bulk material, is the Plank constant, is the reduced mass of the exciton, is radius of the quantum dots, is the charge of electron. is the permittivity of the free space, and is dielectric constant of the material. In this approximation, it is assumed that the particles are spherical in nature, and the motions of electron and hole can be well described in terms of their effective masses. The effect of surface atoms are also ignored in this approximation. For MoS2, the reduced mass of exciton is 0.27 times that of the free electron, the bulk bandgap is 1.29 eV, and the permittivity is about 6.8 times that of free space. The calculated value of the bandgap for MoS2 quantum dots is found to be 1.9 eV. Similarly for WS2 the bandgap was found to be 2.7 eV. In case of WS2, the reduced mass was taken as 0.146 times that of the free electron mass, the dielectric constsnt of the material was taken as 4.13 while theoretically calculation of the band gap. In the Equation 1, the second and third terms of the equation are the kinetic energy and electrostatic attraction terms, respectively. It is the kinetic energy term which dominates in general, and results a blue-shift in the bandgap as the size of the quantum dots decreases. As the size of QDs gets smaller the bandgap gets larger because the valence band moves to more positive potential values, and the conduction band moves to more negative potential values. In most of the semiconducting QDs, the increase in the bandgap is due to the conduction band moving to more negative potential values, and the valence band moves very little (considering potential relative to NHE). Gan et al. (2015); Wu et al. (2005); Parsapour et al. (1996) The calculated bandgap is found to be different from the experimentally observed values using Tauc plot which is due to the different size of the QDs, hence there might be variation in the reduced mass of the excitons and the variation in the dielectric constants values at the nanometer scale.
@@
The photo-response of R6G molecules in the absence as well as in the presence of QDs was examined by using fluorescence spectroscopic techniques. The steady state fluorescence spectra of R6G molecules in the absence and presence of MoS2 and WS2 QDs are recorded by fluorimeter using 430 nm excitation wavelength, and emission was recorded in wavelength range of 520-620 nm. The steady state fluorescence spectra of R6G molecules in the absence and presence of MoS2, and WS2 are shown in Fig. 5. The R6G molecules show characteristic emission peak around 550 nm.
@@
In Fig. 5, it is found that in the presence of QDs, the fluorescence (FL) intensity of R6G decreases. The decrease in FL intensity confirms quenching due to QDs. The FL quenching are expected to be dynamic or static in nature, which can be confirmed by time resolved fluorescence study. In order to explore the possibility of charge transfer from R6G to QDs, FL lifetime of R6G molecules were recored in the absence as well as in the presence of the QDs by using time-correlated single photon counting (TCSPC) set-up. The fluorescence lifetime decay spectra of R6G molecules, in the absence and presence of QDs is shown in Fig. 6.
@@
Fig. 6 shows the FL lifetime decay curve of R6G molecules in the presence of MoS2 and WS2 QDs, and in the absence of the QDs. These decay curves were fitted with exponential decaying functions of following form: Patel, Sahoo, and Mohanty (2014)
where is the number of discrete decay components, are the decay times, and represents the weighing parameter associated with the decay. It is observed that in presence of QDs, the fluorescence decay curves are steeper in comparison to pure R6G. The decay curve of R6G in absence of QDs is fitted with single-exponential () decaying function, while in presence of QDs the curves are fitted with bi-exponential () decaying function. The average fluorescence lifetimes of R6G molecules in absence, as well as in presence of the QDs are calculated. The average fluorescence lifetime of pure R6G was found to be 4.0 ns. In presence of different QDs, the FL lifetime decreases for MoS2 QDs the FL lifetime is found to be 1.9 ns. The FL lifetime is found to be 3.8 ns in presence of WS2 QDs. In all cases, in the presence of QDs, there is a significant decrease in the FL lifetime of R6G molecules. In presence of TMDC QDs, the fluorescence intensity as well as the fluorescence lifetime of R6G molecules are found to decrease. The decrease in lifetime indicates that there is some charge/energy transfer from the R6G molecules to QDs. The decrease in the FL lifetime is prominant in case of MoS2 QDs. In order to understand the possible reasons for decrease in the FL lifetime of the R6G molecules in the presence of QDs, absorption spectra of QDs are plotted along with the emission spectrum of R6G molecules, which is shown in Fig. 7.
@@
In Fig. 7, it is shown that the emission spectrum of R6G does not overlap with the absorption spectra of TMDCs QDs. This cancels the possibilities of energy transfer from R6G to QDs, still there is decrease in the FL intensity as well as FL lifetime of the R6G molecules in presence of QDs, which is due to possibility of charge transfer from the photo-excitted R6G molecules to the QDs. In the process of charge transfer, the fluorophores show quenching in the FL intensity, as well as FL lifetimes. Using FL lifetime data, the charge transfer efficiency from R6G molecules to TMDC QDs has been calculated using following equation:
@@
In summary, the MoS2 and WS2 QDs were synthesized using ultrasonication method. The photo-induced charge transfer between fluorescent R6G molecules and TMDC QDs of MoS2 and WS2 was studied using fluoresce spectroscopic techniques. The QDs of MoS2 and WS2 were used as acceptors for charge transfer, which could be further explored to developing QDs-molecules based optoelectronic devices. Maximum charge transfer efficiency was observed in case of MoS2 QDs which was due to a wide energy difference between the LUMO of the R6G molecules and the conduction energy level of the MoS2 QDs, it had provided a thermodynamically favourable condition of the charge transfer.

### v2

@@ Abstract
In this paper, we have explored the charge transfer mechanism in transition metal dichalcogenide (TMDC) quantum dots (QDs) of molybdenum disulfide () and tungsten disulfide (). Rhodamine 6G (R6G), a dye from the rhodamine family, has been employed as the fluorescent molecule, with MoS2 and WS2 QDs acting as electron acceptors in the photo-induced charge transfer process. The TMDC QDs were synthesized using a top-down approach and characterized through transmission electron microscopy (TEM), UV-Vis spectrophotometry, and fluorimetry. TEM images revealed well-dispersed particles measuring 2 nm in size. These QDs exhibit strong fluorescence emission when excited with light at wavelengths below 350 nm. Under light exposure, photons generate charges in the fluorescent dye molecules, and the TMDC QDs facilitate the charge transfer process. The charge transfer phenomenon was investigated using time-correlated single photon counting (TCSPC), a time-resolved fluorescence spectroscopic technique. The time-resolved fluorescence study indicated a change in the fluorescence (FL) lifetime of R6G molecules in the presence of QDs. The FL lifetime of R6G molecules without QDs was found to be 4.0 ns, which decreased to 1.9 ns and 3.8 ns in the presence of MoS2 and WS2 QDs, respectively. This reduction in FL lifetime suggests that the MoS2 and WS2 QDs provide an additional pathway for photo-generated electrons in the excited state of R6G molecules. This research can be extended to optoelectronic devices, where charge transfer is crucial for device efficiency and performance.
@@
In recent years, different kinds of quantum dots (QDs) have been extensively researched because of their numerous applications in cell imaging, photovoltaic devices, light-emitting devices, and more. Xu et al. (2021); Li et al. (2024); Wang et al. (2020); Sadeghi et al. (2019); Li et al. (2016); Xue et al. (2019); Meng et al. (2024) The most appealing characteristic of these materials is their size-dependent opto-electronic properties, making them ideal candidates for designing various opto-electronic devices.Li et al. (2015); Litvin et al. (2017) The fascinating properties of QDs have been investigated in composite systems or devices through their combination with various materials, including molecules, other QDs, and 2D nanomaterials. When QDs are paired with biomolecules, they are utilized for bio-imaging, while QDs in a core-shell configuration are explored for optoelectronic devices. These QDs, when sensitized with other semiconductors, have been used to create cost-effective, next-generation photovoltaic devices for various applications. QDs are favored in these applications due to their size-dependent optical and electronic properties. The interaction between QDs and their surrounding media or materials is critical for device performance. Although the electronic interactions between organic molecules and fluorescent QDs have been well studied, systems where QDs are coupled with other inorganic molecules have been less explored. These interactions differ from those in QD-molecular systems because inorganic materials have a continuum of electronic states, whereas molecular acceptors have discrete electronic states. When QDs are coupled with other inorganic materials, they can donate electrons, as observed in photovoltaic devices, or act as electron donors/acceptors, as seen in light-emitting diodes (LEDs). In these applications, electron transfer involves a specific direction of transfer. Therefore, understanding the factors driving electron transfer in these composite systems is crucial for better comprehending and further exploiting the unique properties of QDs in optoelectronic devices.Algar and Krull (2008); Wang et al. (2021); Chakraborti et al. (2016); Zhu, Zhang, and Xia (2018); Lima et al. (2022); Wu et al. (2020); Rani et al. (2020)
@@
There are various studies on different kind of QDs, where these QDs have been used as acceptor/donor or both of the electrons. They have also been explored for donor or acceptor in the resonance energy transfer. In recent years researchers have explored various kinds of layered based QDs in place of traditional QDs, e.g., CdS, CdSe, PbS, etc.Kharangarh et al. (2023); Sariga et al. (2023); Gan et al. (2015); Guo et al. (2019) In optoelectronic applications, QDs function as light-harvesting and charge-separation centers. A crucial step in this process is the transfer of electrons or holes from the photoexcited QDs to nearby acceptors through a nanoscale interface. The overall efficiency of the device relies on these charge transfer processes, which are influenced by competing loss pathways. Therefore, understanding and controlling interfacial charge transfer processes is vital for improving QD-based devices. The size of the QDs also significantly impacts the charge transfer process, with studies showing that the electron transfer rate increases as the QD size decreases. Zhu et al. (2016)
@@
In recent years, various 2D layered nanomaterials have been transformed into QDs, which exhibit significantly different properties compared to their 2D structures and bulk counterparts. Specifically, 2D layered transition metal dichalcogenides (TMDCs) have been converted into QDs using various methods. Among these TMDC-based QDs, MoS2 and WS2 QDs have been extensively studied. They exhibit intriguing optical and electronic properties, with their electronic structure being tunable by adjusting their size and surface morphology. These QDs have been applied in various fields, including sensors, LEDs, photodetectors, and cell imaging. Yan et al. (2016); Ghorai et al. (2017); Singh et al. (2019); Guo and Li (2020)
@@
Several studies have used TMDC QDs as energy donors in hybrid systems due to their photoluminescence in the UV region. There is extensive research on energy transfer and charge transfer in various nanomaterials and their applications in sensors, photodetectors, and solar cells. In this work, we have investigated charge transfer in two different types of fluorescent materials: fluorescent dye molecules and QDs. Typically, QDs are used as donors or acceptors, and dye molecules as acceptors or donors in energy transfer studies, which often focus on the energy transfer and its applications. Here, we have explored the potential for charge transfer from dye molecules to QDs, providing valuable insights for developing optoelectronic devices based on fluorescent nanomaterials. In this paper, we used a common dye molecule, R6G, as the source of photo-excited electrons and TMDC QDs as the electron acceptors.
@@
The TMDC QDs were synthesized using top-down strategy as mentioned in Long et al. with slight modification.Long et al. (2016) In typical synthesis procedure, 50 mg of TMDC powder was dissolved in 50 mL of NMP solution with constant stirring for 15 hours at 29 ∘C. The resulting dark suspension was centrifuged at 10,000 rpm for 40 min. The supernatant of the solution was collected and used for further studies. The schematic diagram of the synthesis of TMDC quantum dots is shown in Fig 1.
@@
In our study, prior to exploring charge transfer mechanism, the quantum dots (QDs) were characterized using transmission electron microscopy (TEM), fluorescence spectroscopy, and absorption spectroscopy. TEM images were captured with a JEOL-2010 TEM operating at 200 kV. The absorption spectra of the QDs were obtained using a Shimadzu (UV-2450) UV-Vis spectrophotometer. In order to study charge transfer in these TMDCs QDS, the QDs were mixed with a fixed quantity of R6G solution. For charge transfer studies, 0.5 mL aqueous solution of 10-6 M R6G was further diluted to 3.5 mL, and 1 mL of QDs solution was mixed to R6G solution. In all the samples the NMP and water contents were kept same. The steady-state fluorescence spectra of R6G molecules were captured both in the absence and presence of TMDCs QDs using a fluorimeter with an excitation wavelength of 430 nm. The time-resolved fluorescence analysis was conducted using a time-correlated single-photon counting module (TCSPC) (FL920, Edinburgh Instruments, UK), with excitation provided by a diode laser at 465 nm.
@@
The absorption spectra of MoS2 and WS2 quantum dots are shown in Fig 2. A clear absorption peak around 325 nm is observed, which is a characteristic peak of WS2 quantum dots and it is due to excitonic features of the QDs. Similarly, for MoS2 QDs a band edge around 380 nm is observed. The color of suspension and the absorption peaks are similar to the previous reported results. Coleman et al. (2011); Long et al. (2016) When the MoS2 and WS2 are in the bulk or nanosheet form, they exhibit quite different absorption patterns having two distinct peaks around 600 nm which are known as and excitonic peaks. While in case of QDs these and excitonic peaks disappear and absorption is observed in UV region. The pronounced blue shift observed in the absorption spectra of the synthesized MoS2 and WS2 QDs can be attributed to the quantum confinement effect and edge effects that become significant as the lateral dimensions of the QDs are reduced to a few tens of nanometers. Li et al. (2017); Liu et al. (2014) In order to have insight of the band gaps of these QDs, the Tauc plots of these QDs are drawn. The Tauc plots of the quantum dots are shown in Fig 3.
@@
In case of semiconductors, it has been observed the bandgap increases with decrease in the size of the semiconducting quantum dots. The size dependent bandgap is given by following equation:Wu et al. (2005)
@@
In this expression, represents the bandgap of the bulk material, is Planck’s constant, is the reduced mass of the exciton, is the quantum dot radius, is the electron charge, is the permittivity of free space, and is the dielectric constant of the material. This approximation assumes the particles are spherical, and the motions of the electron and hole are described by their effective masses. The effect of surface atoms are also ignored in this approximation. For MoS2, the reduced mass of exciton is 0.27 times that of the free electron, the bulk bandgap is 1.29 eV, and the permittivity is about 6.8 times that of free space. The calculated value of the bandgap for MoS2 quantum dots is found to be 1.9 eV. Similarly for WS2 the bandgap was found to be 2.7 eV. In case of WS2, the reduced mass was taken as 0.146 times that of the free electron mass, the dielectric constant of the material was taken as 4.13 while theoretically calculation of the band gap. In the Equation 1, the second and third terms of the equation are the kinetic energy and electrostatic attraction terms, respectively. It is the kinetic energy term which dominates in general, and results a blue-shift in the bandgap as the size of the quantum dots decreases. As the size of QDs decreases, the bandgap increases because the valence band shifts to more positive potentials, while the conduction band shifts to more negative potentials. In most semiconducting QDs, this increase in bandgap is primarily due to the conduction band moving to more negative potentials, with only a slight shift in the valence band. (considering potential relative to NHE). Gan et al. (2015); Wu et al. (2005); Parsapour et al. (1996) The calculated bandgap is found to be different from the experimentally observed values using Tauc plot which is due to the different size of the QDs, hence there might be variation in the reduced mass of the excitons and the variation in the dielectric constants values at the nanometer scale.
@@
The photo-response of R6G molecules, both in the absence and presence of QDs, was investigated using fluorescence spectroscopy. The steady-state fluorescence spectra of R6G molecules, with and without MoS2 and WS2 QDs, were recorded using a fluorimeter with a 430 nm excitation wavelength. Emission measurements were taken across a wavelength range of 520-620 nm. The steady state fluorescence spectra of R6G molecules in the absence and presence of MoS2, and WS2 are shown in Fig. 6. The R6G molecules show characteristic emission peak around 550 nm.
@@
In Fig. 6, it was observed that the fluorescence intensity of R6G decreases in the presence of QDs, indicating quenching by the QDs. This quenching could be either dynamic or static, which can be confirmed through time-resolved fluorescence studies. To investigate potential charge transfer from R6G to the QDs, the fluorescence lifetime of R6G molecules was measured both with and without the QDs using a time-correlated single photon counting (TCSPC) setup. The resulting fluorescence lifetime decay spectra of R6G in both scenarios are depicted in the Fig. 7.
@@
Fig. 7 presents the fluorescence lifetime decay curves of R6G molecules both in the presence of MoS2 and WS2 quantum dots (QDs) and in their absence. These decay curves were analyzed using exponential decay functions of the following form: Patel, Sahoo, and Mohanty (2014)
The expression for the fluorescence decay includes as the number of discrete decay components, with representing the decay times, and as the weighting factors corresponding to the decay component. It was observed that the fluorescence decay curves steepened in the presence of quantum dots (QDs) compared to pure R6G. While the R6G decay curve alone was well-fitted with a single-exponential function (), in the presence of QDs, a bi-exponential function () was required for fitting. The average fluorescence lifetime for R6G alone was 4.0 ns, but this decreased to 1.9 ns with MoS2 QDs and 3.8 ns with WS2 QDs. The reduction in fluorescence lifetime suggests charge or energy transfer from R6G molecules to the QDs, with the effect being more pronounced for MoS2 QDs. To further investigate the reasons for this decrease in fluorescence lifetime, the absorption spectra of the QDs were compared with the emission spectrum of R6G molecules, as depicted in Fig. 8.
@@
In Fig. 8, the emission spectrum of R6G does not overlap with the absorption spectra of TMDC QDs, which rules out energy transfer from R6G to the QDs as a cause for the observed reduction in fluorescence intensity and lifetime. Instead, the decrease is likely due to charge transfer from photo-excited R6G molecules to the QDs. This charge transfer process leads to quenching of both fluorescence intensity and lifetime. The charge transfer efficiency from R6G to TMDC QDs was calculated using the following equation:
@@
In summary, the MoS2 and WS2 QDs were synthesized using chemical assisted exfoliation method. The photo-induced charge transfer between fluorescent R6G molecules and TMDC QDs of MoS2 and WS2 was studied using fluorescence spectroscopic techniques. The QDs of MoS2 and WS2 were used as acceptors for charge transfer, which could be further explored to developing QDs-molecules based optoelectronic devices. Maximum charge transfer efficiency was observed in case of MoS2 QDs which was due to a wide energy difference between the LUMO of the R6G molecules and the conduction energy level of the MoS2 QDs, it had provided a thermodynamically favourable condition of the charge transfer.